the perimeter of a triangle (the sum of its three sides) is 145.one of the sides is twice the length of the smallest side, and the third is five less than three times the smaller side.Find the lenght of each side
Answer by stanbon(75887)

the perimeter of a triangle (the sum of its three sides) is 145.one of the sides is twice the length of the smallest side, and the third is five less than three times the smaller side.
Find the length of each side
-------------------
a + b + c = 145
b = 2a
c = 3a-5
-----
a + 2a + 3a-5 = 145
6a = 150
------------------------
a = 25
b = 2a = 50
c = 3a-5 = 70
====================
Cheers,
Stan H.
===========